OpenStudy (anonymous):

One solution, infinite, or no solution 22. Y=x+3 y=x-1 23. y=2x-1 3y=6x-5 24.3x+y=2 4y=12-12x 26.y=2x-2 2y=4x-4 27. y-x=5 3y=3x+15 28. 2x+2y=4 12-3x=3y

OpenStudy (anonymous):

Please help

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

so for 22 would it be no solution because earlier the lines didn't intersect?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Thanks

OpenStudy (loser66):

first equation y = 2x -2 divided by 2 the 2nd one y = 2x -2 again , let them equal, so 2x -2 =2x -2 or -2 =-2 , bingo, it's always right for any x. so, the conclusion is infinite solution, whatever x is, I always have -2 = -2 . Make sense to you?